# MMSci: A Dataset for Graduate-Level Multi-Discipline Multimodal Scientific Understanding

### v2

@@ Abstract
The rapid development of Multimodal Large Language Models (MLLMs) is making AI-driven scientific assistants increasingly feasible, with interpreting scientific figures being a crucial task. However, existing datasets and benchmarks focus mainly on basic charts and limited science subjects, lacking comprehensive evaluations. To address this, we curated a multimodal, multidisciplinary dataset from peer-reviewed, open-access Nature Communications articles, spanning 72 scientific disciplines. This dataset includes figures such as schematic diagrams, simulated images, macroscopic/microscopic photos, and experimental visualizations (e.g., western blots), which often require graduate-level, discipline-specific expertise to interpret. We developed benchmarks for scientific figure captioning and multiple-choice questions, evaluating six proprietary and over ten open-source models across varied settings. The results highlight the high difficulty of these tasks and the significant performance gap among models. While many open-source models performed at chance level on the multiple-choice task, some matched the performance of proprietary models. However, the gap was more pronounced in the captioning task. Our dataset also provide valuable resource for training. Fine-tuning the Qwen2-VL-2B model with our task-specific multimodal training data improved its multiple-choice accuracy to a level comparable to GPT-4o, though captioning remains challenging. Continuous pre-training of MLLMs using our interleaved article and figure data enhanced their material generation capabilities, demonstrating potential for integrating scientific knowledge. The dataset and benchmarks will be released to support further research.
Figure 1: The top 30 out of 72 science subjects [^2] https://www.nature.com/nature/browse-subjectswith the most articles in our dataset MMSci. The corresponding numbers of figures are shown in brackets.
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) (Li et al., 2023; Zhu et al., 2023; Liu et al., 2024; Chen et al., 2024b; Bai et al., 2023b; Achiam et al., 2023; Team et al., 2023; Anthropic, 2024a; Wang et al., 2024a), have demonstrated remarkable capabilities in solving problems involving visual context. These range from everyday scenes (Antol et al., 2015), to reading documents (Tito et al., 2021), understanding diagrams (Kembhavi et al., 2016), interpreting charts (Kahou et al., 2017; Masry et al., 2022), solving math (Lu et al., 2023), and multi-discipline problems (Yue et al., 2023; 2024). The growing capabilities and intelligence of MLLMs make them promising as AI-driven scientific assistants capable of solving problems and assisting in research in various scientific domains. A critical aspect of this assistance is interpreting the figures in research articles, which often contain and convey rich, compressed, and complex information, requiring expert knowledge in specific scientific fields for proper understanding.
@@
However, the ability of current MLLMs to comprehend diverse complex scientific figures across diverse disciplines remains largely unexplored. Existing research has primarily focused on chart interpretation (Kahou et al., 2017; Masry et al., 2022; Roberts et al., 2024; Wang et al., 2024b), which requires minimal domain-specific knowledge. In contrast, figures in scientific articles are far more varied, including microscopy and spectroscopy images, astronomical images, maps, 3D models, molecular structures, geological models, phylogenetic trees, electropherograms, waveforms, heatmaps, spectrograms, etc. Interpreting these figures often requires expert, typically graduate-level, knowledge in specific domains. Furthermore, datasets addressing figures beyond charts tend to focus on a limited range of subjects, lacking coverage of the full spectrum of scientific fields Li et al. (2024).
To bridge the gap, we collected a multimodal, multi-discipline dataset MMSci from high-quality, open-access articles published in Nature Communications [^3] https://www.nature.com/ncomms/, which are freely and permanently available upon publication under a Creative Commons Attribution 4.0 International (CC BY) license [^4] More details can be found at https://www.nature.com/ncomms/open-access. This dataset spans 72 scientific disciplines, primarily within the natural sciences (the top 30 subjects with most articles can be seen in Figure 1). It includes over 131k articles and 742k figures, featuring a wide range of figures types in these research articles (seven major image types are shown in Figure 2). These figures often require graduate-level, domain-specific expertise to interpret, going beyond basic chart comprehension. To evaluate MLLMs’ understanding of complex multimodal scientific figures, we developed a benchmark with tasks including scientific figure captioning and multiple-choice questions, offering a thorough assessment of models’ ability to analyze diverse and advanced scientific figures and content.
@@
Our evaluation revealed challenges across tasks and notable performance gaps among current MLLMs. For the multiple-choice questions, many open-source models performed no better than random guessing. However, some models, such as Qwen2-VL-7B (Wang et al., 2024a) and MiniCPM-V-2.6 Yao et al. (2024), showed strong performance, comparable to some proprietary models like Gemini-1.5-Flash (Reid et al., 2024), and Claude-3-Opus (Anthropic, 2024a). GPT-4o (Achiam et al., 2023) and Claude-3.5-Sonnet (Anthropic, 2024b) are the leading models, outperformed even computer science master’s students, underscoring their potential as scientific assistants. This also highlights the task’s difficulty and the demand for domain-specific knowledge. All models struggle to generate precise captions, particularly with nuanced semantics. However, GPT-4V and GPT-4o perform significantly better than other models when grounded on article abstracts.
@@
Additionally, our dataset provides a vast collection of high-quality research articles and figures across diverse subjects, which can be leveraged as training resources to enhance MLLMs’ understanding of multimodal scientific content. We experimented with constructing visual supervised fine-tuning data, including the task-specific data converted into single- and multi-turn conversations. This data significantly improved the Qwen2-VL-2B model (Wang et al., 2024a), achieving the highest overall multiple-choice accuracy on our benchmark, though improving captioning performance remained challenging. Furthermore, we pre-trained MLLMs on interleaved article text and figure images, which led to improved performance in material generation, a downstream task in material sciences.
@@
Overall, our contributions are threefold: (1) Data diversity, scope and quality: Our dataset is uniquely composed of high-quality, peer-reviewed academic articles covering 72 diverse scientific disciplines, featuring a wide range of figure types beyond charts. (2) Challenging benchmark: Our benchmark includes tasks with diverse settings to ensure a comprehensive assessment. The evaluation highlights the challenges of the task and the limitations of current MLLMs in effectively interpreting figures from scientific literature. (3) Rich training resources: Our dataset provides a valuable training resource. We created task-specific multimodal fine-tuning data and interleaved article and figure data for continuous MLLM pre-training. Our findings highlight the potential of this dataset to improve models’ comprehension of scientific knowledge.
Table 1: Comparison with prior scientific figure understanding benchmark datasets.
@@
Benchmark Dataset Data Source Peer-reviewed # Subjects Image Type Annotations
@@ 2 Related Dataset Work
Scientific Figure Understanding
Scientific figures in academic articles convey rich, valuable information, and there has been extensive research on evaluating the understanding of these figures. As seen in Table 1, existing datasets primarily focus on relatively simple chart figures, which require minimal scientific knowledge but general chart interpretation capacities. Early efforts targeted data visualization figures, such as synthetic datasets of plots and charts (Chen et al., 2020; Kahou et al., 2017; Kafle et al., 2018). To capture more diverse and complex chart figures, FigureSeer (Siegel et al., 2016) and SciCap (Yang et al., 2023) extracted figures from computer science (CS) papers on arXiv. SciFiBench (Roberts et al., 2024) expanded on SciCap’s chart figures by introducing figure-to-caption and caption-to-figure matching tasks, while CharXiv (Wang et al., 2024b) hand-picked chart figures from arXiv papers. These datasets focus exclusively on chart figures. ArxivQA/Cap (Li et al., 2024) extended the scope by collecting papers from 32 subjects on arXiv, including open-category image types beyond charts. However, the collection still heavily focuses on CS and math, with less comprehensive coverage of the natural sciences. Moreover, since arXiv papers are not peer-reviewed, their quality is not guaranteed. In contrast, our dataset includes peer-reviewed articles from Nature Communications, spanning 72 disciplines and covering a wide range of natural science subjects.
@@
The figures are highly diverse and typically require graduate-level expertise in specific subjects to interpret, ensuring the dataset’s quality, diversity, and complexity.
Multimodal Science Problems
With the advances of MLLMs, many studies have focused on evaluating their ability to solve scientific problems involving visual context. However, these datasets typically emphasize models’ ability to “read” and “see” simple image content for use in solving the problem, rather than testing their “understanding” of complex scientific figures. The images in these datasets are relatively simple and usually do not require expert scientific knowledge for interpretation. For example, ScienceQA (Lu et al., 2022) focuses on K1-12 level problems, while SciBench (Wang et al., 2023) is limited to three disciplines: physics, chemistry, and mathematics. MMMU (Yue et al., 2023) covers subjects such as art, business, history, health, humanities, and technology, but its coverage of natural science subjects is limited, and understanding images is not the primary challenge. Our work, in contrast, with the focus on the understanding of complex scientific figures that require graduate-level, domain-specific knowledge across scientific disciplines. It can also be potentially used for constructing multimodal science problems, which we leave for future exploration.
@@ 3 Data Curation
Our dataset was collected from the Nature Communications website, comprising open-access, peer-reviewed papers across five major categories and 72 subjects. The top 30 subjects are shown in Figure 1, with the full list of all 72 subjects provided in the Appendix, Table 6. Various information regarding each article is easily accessible on this website, providing a user-friendly platform for obtaining all necessary data. For each article, we collected information including the title, abstract, main body content, and references, directly from their respective sections on the article’s webpage (e.g., https://www.nature.com/articles/xxx, where “xxx” is the article’s unique ID). Figures and their captions were obtained from a dedicated figures page under the article’s homepage (e.g., https://www.nature.com/articles/xxx/figures), eliminating the need to extract figures from PDF files and thus ensuring image quality. We used pylatexenc to convert LaTeX expressions of mathematical formulas in the article text and figure captions into plain text. [^5] https://github.com/phfaist/pylatexenc Since these papers are all peer-reviewed and the text, figures, and captions are readily available from the website, ensuring the data is both authentic and high-quality. We thus did not perform additional filtering or content extraction. We crawled articles up to the date of 2024/04/15. The resulting source dataset comprises 131,393 articles and 742,273 figures. More statistics are shown in Table 2.
Sub-caption Extraction
@@
Many figures in the dataset consist of multiple sub-figures in a single image, with captions that include a main caption and descriptions of each sub-figure (sub-caption), as illustrated in Figure 3. We developed a regular expression matching function to identify sub-figure indices at the beginning of sentences in alphabetical order (a to z), extracting and identifying 514,054 sub-captions/figures, which aids in the consecutive construction of our benchmark.
@@
Figure 3: Illustration of the benchmark data in MMSci. This example is taken from (Guo et al., 2024). The left side shows a figure with multiple sub-figures, accompanied by a main caption (bolded) and sub-captions (highlighted in different colors), each corresponding to a sub-figure. These sub-figures and sub-captions are used to construct tasks for figure captioning (upper right), sub-figure to sub-caption matching (center right), and sub-caption to sub-figure matching (lower right).
Heterogeneous Figure Types in MMSci
@@
We categorized the types of (sub-)figures in MMSci into seven major categories based on a subset of the figures, focusing on the smallest individual components, such as sub-figures when present. Following this manual review, we used GPT-4o to classify the images within the benchmark test set (see benchmark data splits in the next section). Examples of image types are shown in Figure 2, with detailed statistics provided in the Appendix (Section A.1.3). In addition to charts in previous benchmarks, which make up half of the figures, we identified six other major types that vary significantly across different subjects.
@@ 4 Benchmarks
We developed two benchmark tasks MMSciCap and MMSciQA with varying settings to comprehensively test models’ understanding of scientific figures and content, as shown in Figure 3.
MMSciCap: Scientific Figure Captioning
@@
Scientific figure captioning in MMSci presents unique challenges compared to typical natural image captioning. Unlike natural images, interpreting figures in Nature Communications articles often requires graduate-level expertise in specific subjects and grounding in the article’s content. Additionally, these captions are significantly more detailed. In MMSci, captions average 153 words, much longer than those for natural images and ArxivCap (Li et al., 2024). This complexity makes scientific figure captioning in our benchmark more demanding. In our task, we evaluate the scientific figure captioning with two settings: (1) ungrounded figure captioning: The model generates captions without access to any additional article content. (2) abstract-grounded figure captioning: The model is provided with the paper abstract to offer context for the figure. For evaluation metrics, we consider overlap-based metrics ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and the similarity-based metric BERTScore (Zhang et al., 2019), which compare the generated captions to the reference captions.
@@
Additionally, we use a Large Language Model (LLM)-based metric, modified from FActScore (Min et al., 2023) for evaluating detailed and complex scientific figure captions in our task. The key idea is to use LLMs to break down the generated caption $y$ into a set of atomic units, denoted as $\mathcal{A}_{y}$. Each atomic unit represents an independent description of either the overall figure or individual sub-figures. We then evaluate whether each atomic unit is supported by the ground-truth caption $\mathcal{C}$. To achieve a more fine-grained evaluation, the LLM assigns a score $\phi(a,\mathcal{C})$ to each atomic unit $a\in\mathcal{A}_{y}$ on a scale from 0 to 1, representing the degree of support from the ground-truth caption. Additionally, a brevity penalty is used to account for overly concise captions. The overall formulation is defined as follows:
$\begin{gathered}f(y)=\frac{1}{|\mathcal{A}_{y}|}\sum_{a\in\mathcal{A}_{y}}\phi(a,\mathcal{C})\cdot\text{exp}(\text{min}(1-\frac{\gamma}{\mathcal{A}_{y}},0)).\end{gathered}$
@@
We set $\gamma$ to 10 in our evaluation. Note that this metric focuses on precision rather than recall. We also use G-Eval (Liu et al., 2023b) to compare the generated caption with the reference caption on a scale of 1 to 5. We use GPT-4o for the LLM-based evaluations.
MMSciQA: Figure Caption Matching
@@
We also construct multiple-choice questions to evaluate models’ ability to understand (sub-)figures and match them with the correct captions. To comprehensively assess different capabilities and scenarios, we design three settings: (1) Figure-to-Caption (Fig2Cap): The options include the correct main caption for a whole figure and three other main captions from different figures within the same article. This tests the model’s ability to summarize the entire figure, demonstrating a holistic understanding of the entire figure. (2) Subfigure-to-Subcaption (SubFig2Cap): This setting evaluates the model’s ability to locate a sub-figure and match it with its corresponding sub-caption. A random sub-figure is selected, along with its correct sub-caption and three other sub-captions within the figure as choices. This challenges the model to first identify the correct sub-figure, then interpret the captions to select the right answer. (3) Subcaption-to-Subfigure (SubCap2Fig): In this reverse setting, the model is given a sub-caption and must select the matching sub-figure from all available sub-figures in the figure. This requires the model to examine each sub-figure and compare it with the captions. All sub-figures in the figure are included as choices.
Data Split
To manage evaluation costs and align with popular benchmark sizes, we allocated 1% of articles from each subject to the test set and another 1% to the validation (dev) set, resulting in 1,418 test articles and 1,414 validation articles, with 5 to 50 articles per subject. Each test sample was derived from a single article, ensuring no content reuse. For the captioning task, captions were required to exceed 50 words. Each task and setting ultimately included around 1,200 samples, balancing coverage, diversity, and cost for effective benchmarking.

### v3

@@ Abstract
Scientific figure interpretation is a crucial capability for AI-driven scientific assistants built on advanced Large Vision Language Models. However, current datasets and benchmarks primarily focus on simple charts or other relatively straightforward figures from limited science domains. To address this gap, we present a comprehensive dataset compiled from peer-reviewed Nature Communications articles covering 72 scientific fields, encompassing complex visualizations such as schematic diagrams, microscopic images, and experimental data which require graduate-level expertise to interpret. We evaluated 19 proprietary and open-source models on two benchmark tasks, figure captioning and multiple-choice, and conducted human expert annotation. Our analysis revealed significant task challenges and performance gaps among models. Beyond serving as a benchmark, this dataset serves as a valuable resource for large-scale training. Fine-tuning Qwen2-VL-7B with our task-specific data achieved better performance than GPT-4o and even human experts in multiple-choice evaluations. Furthermore, continuous pre-training on our interleaved article and figure data substantially enhanced the model’s downstream task performance in materials science. We will release our dataset to support further research.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: The top 20 out of 72 science subjects with the most articles in our dataset MMSci. The corresponding numbers of papers and figures (in brackets) are shown.
@@
Recent advancements in Large Vision Language Models (LVLMs) (Li et al., 2023; Zhu et al., 2023; Liu et al., 2024; Chen et al., 2024c; Bai et al., 2023b; Achiam et al., 2023; Team et al., 2023; Anthropic, 2024a; Wang et al., 2024a), have demonstrated remarkable capabilities in solving problems involving visual context. The growing capabilities of LVLMs make them promising as AI-driven scientific assistants capable of solving problems and assisting in research in various science domains. A critical aspect of this assistance is interpreting the figures in research articles, which often contain rich, compressed, and complex information, requiring domain-specific expertise to understand.
@@
Current evaluations of LVLMs on scientific figures focus primarily on bar chart interpretation (Kahou et al., 2017; Masry et al., 2022; Roberts et al., 2024; Wang et al., 2024b), and relatively easy figures within limited science domains (Yue et al., 2023, 2024; Li et al., 2024; Chen et al., 2024a). However, figures in scientific articles are far more varied, including microscopy and spectroscopy images, astronomical images, maps, 3D models, molecular structures, geological models, phylogenetic trees, electropherograms, waveforms, heatmaps, spectrograms, etc. Interpreting these figures often requires expert, typically graduate-level, knowledge in specific domains.
To bridge this gap, we introduce MMSci, a comprehensive multimodal dataset curated from open-access Nature Communications articles [^1] https://www.nature.com/ncomms/ under CC BY 4.0 license [^2] https://www.nature.com/ncomms/open-access. The dataset encompasses 72 scientific disciplines, containing 131k articles and 742k figures across diverse visualization types (Figure 2), with discipline distribution shown in Figure 1 (only shows top-20 due to space constraints). To evaluate LVLMs’ comprehension of these complex scientific figures requiring graduate-level expertise, we developed benchmark tasks for figure captioning and multiple-choice questions across different settings.
@@
Our evaluation revealed significant performance gaps among current LVLMs across tasks. For multiple-choice questions, many open-source models performed no better than random guessing. However, some models, such as Qwen2-VL-7B (Wang et al., 2024a) and MiniCPM-V-2.6 (Yao et al., 2024), demonstrated strong performance comparable to proprietary models like Gemini-1.5-Flash (Reid et al., 2024) and Claude-3-Opus (Anthropic, 2024a). GPT-4o (Achiam et al., 2023) and Claude-3.5-Sonnet (Anthropic, 2024b) emerged as the leading models, significantly outperforming other evaluated models. We also conducted human expert evaluations. The results revealed that the leading models achieved performance comparable to or exceeding domain experts, demonstrating their potential as cross-domain scientific assistants. This performance also highlights our tasks’ difficulty and the importance of domain-specific knowledge. While all models struggled with generating precise figure captions, particularly for nuanced semantics, Claude-3.5-Sonnet and GPT-4o still showed markedly improved performance.
@@
Additionally, our dataset provides a vast collection of high-quality research articles and figures across diverse subjects, which can be leveraged as training resources to enhance LVLMs’ understanding of multimodal scientific content. We experimented with constructing visual supervised fine-tuning data, including the task-specific data converted into instruction-following data. This data significantly improved the Qwen2-VL-7B model (Wang et al., 2024a), achieving the highest overall multiple-choice accuracy on our benchmark, though improving captioning performance remained challenging. Furthermore, we pre-trained LVLMs on interleaved article text and figure images, which led to improved performance in material generation, a downstream task in material sciences.
@@
Overall, our contributions are threefold: (1) Data diversity, scope and quality: Our dataset is uniquely composed of high-quality, peer-reviewed academic articles covering 72 diverse scientific disciplines, featuring a wide range of figure types beyond charts. (2) Challenging benchmark: Our benchmark includes tasks with diverse settings to ensure a comprehensive assessment. The evaluation of models and human experts highlights the challenges of the task. (3) Rich training resources: Our dataset provides a valuable training resource. We created task-specific multimodal fine-tuning data and interleaved article and figure data for continuous LVLM pre-training. Our findings highlight the potential of this dataset to improve models’ comprehension of scientific knowledge.
Table 1: Comparison with prior scientific figure understanding benchmark datasets. *The number of subjects in each work is taken from the original paper that uses different taxonomies, offering a sense of the relative coverage in each work.
@@
Benchmark Dataset Data Source Peer-reviewed # Subjects* Image Type Annotations Training Set
@@ 2 Related Dataset Work
Scientific Figure Understanding.  Scientific figures in academic articles convey rich, valuable information, and there has been extensive research on evaluating their interpretation. As shown in Table 1, existing datasets primarily focus on relatively simple chart figures, which require general chart interpretation skills rather than deep scientific knowledge. Early efforts targeted data visualization figures through synthetic datasets of plots and charts (Chen et al., 2020; Kahou et al., 2017; Kafle et al., 2018). To capture more diverse and complex chart figures, FigureSeer (Siegel et al., 2016) and SciCap (Yang et al., 2023) extracted figures from computer science (CS) papers on arXiv. SciFiBench (Roberts et al., 2024) expanded on SciCap’s chart figures by introducing figure-to-caption and caption-to-figure matching tasks, while CharXiv (Wang et al., 2024b) hand-picked chart figures from arXiv papers. While these datasets focus exclusively on chart figures, ArxivQA/Cap (Li et al., 2024) extended the scope by collecting papers from 32 subjects on arXiv, including various image types beyond charts. However, the collection remains heavily focused on CS and mathematics, with limited coverage of natural sciences. Moreover, since arXiv papers are not peer-reviewed, their quality cannot be guaranteed. In contrast, our dataset comprises peer-reviewed articles from Nature Communications, spanning 72 disciplines and covering a wide range of natural science subjects. We also provide a rich training set for enhancing scientific figure understanding capabilities.
@@
Multimodal Science Problems.  With the advances in LVLMs, recent studies have focused on evaluating their ability to solve scientific problems involving visual context. However, existing datasets primarily assess models’ ability to ”read” and ”see” simple image content rather than testing their ”understanding” of complex scientific figures. The images in these datasets are relatively straightforward and typically do not require expert scientific knowledge for interpretation. For example, ScienceQA (Lu et al., 2022) focuses on K-12 level problems, while SciBench (Wang et al., 2023) is limited to three disciplines: physics, chemistry, and mathematics. MMMU (Yue et al., 2023) and MMMU-Pro (Yue et al., 2024) cover subjects such as art, business, history, health, humanities, and technology, but their coverage of natural science subjects is limited, and image understanding is not the primary challenge. While MMStar (Chen et al., 2024a) includes natural sciences, its coverage is somewhat limited. In contrast, our work focuses on understanding complex scientific figures that require graduate-level, domain-specific knowledge across scientific disciplines. Our dataset can potentially be used for constructing multimodal science problems, which we leave for future exploration.
@@ 3 Data Curation
Source Data Collection.  Our dataset was collected from the Nature Communications website, comprising open-access, peer-reviewed papers across five major categories and 72 subjects. The top 20 subjects are shown in Figure 1, with the full list of all 72 subjects provided in the Appendix, Table 6. Various information regarding each article is easily accessible on this website, providing a user-friendly platform for obtaining all necessary data. For each article, we collected information including the title, abstract, main body content, and references, directly from their respective sections on the article’s webpage (e.g., https://www.nature.com/articles/xxx, where “xxx” is the article’s unique ID). Figures and their captions were obtained from a dedicated figures page under the article’s homepage (e.g., https://www.nature.com/articles/xxx/figures), eliminating the need to extract figures from PDF files and thus ensuring image quality. We used pylatexenc to convert LaTeX expressions of mathematical formulas in the article text and figure captions into plain text. [^3] https://github.com/phfaist/pylatexenc Since these papers are all peer-reviewed and the text, figures, and captions are readily available from the website, ensuring the data is both authentic and high-quality. We thus did not perform additional filtering or content extraction. We crawled articles up to the date of 2024/04/15. The resulting source dataset comprises 131,393 articles and 742,273 figures. More statistics are shown in Table 2.
@@
Sub-caption Extraction.  Many figures in the dataset consist of multiple sub-figures in a single image, with captions that include a main caption and descriptions of each sub-figure (sub-caption), as illustrated in Figure 3. We developed a regular expression matching function to identify sub-figure indices at the beginning of sentences in alphabetical order (a to z), extracting and identifying 514,054 sub-captions/figures, which aids in the consecutive construction of our benchmark.
@@
Figure 3: Illustration of the benchmark data in MMSci. This example is taken from (Guo et al., 2024). The figure (left) contains multiple sub-figures with a main caption (bold) and color-coded sub-captions corresponding to each sub-figure. These sub-figures and sub-captions are used to construct tasks for figure captioning (upper right), sub-figure to sub-caption matching (center right), and sub-caption to sub-figure matching (lower right).
@@
Heterogeneous Figure Types in MMSci We categorized the types of (sub-)figures in MMSci into seven major categories based on a subset of the figures, focusing on the smallest individual components, such as sub-figures when present. Following this manual review, we used GPT-4o to classify the images within the benchmark test set (see benchmark data splits in the next section). Examples of image types are shown in Figure 2, with detailed statistics provided in the Appendix (Section A.1.3). In addition to charts in previous benchmarks, which make up half of the figures, we identified six other major types that vary significantly across different subjects.
@@ 4 Benchmarks
We developed two benchmark tasks with varying settings to comprehensively test models’ understanding of scientific figures and content, as shown in Figure 3.
@@
MMSciCap: Scientific Figure Captioning.  Scientific figure captioning in MMSci poses unique challenges compared to natural image captioning. Interpreting figures from Nature Communications articles requires graduate-level domain expertise and understanding of the article’s context. Moreover, these captions are substantially more detailed, averaging 153 words—significantly longer than those in natural image datasets and ArxivCap (Li et al., 2024). This complexity makes our benchmark particularly challenging. We evaluate scientific figure captioning under three settings: (1) Figure-only captioning: Models generate captions solely from the figure without additional context. (2) Abstract-grounded captioning: Models receive both the figure and the paper’s abstract as context. We also evaluated models’ performance when provided with full article content, limiting this assessment to long context proprietary models due to length constraints, detailed in Appendix A.2.2.
@@
For evaluation metrics, we consider overlap-based metrics BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), the similarity-based metric BERTScore (Zhang et al., 2019), which compare the generated captions to the reference captions, and also the captioning-specific metric CIDEr (Vedantam et al., 2015). Additionally, we use two LLM-based metrics tailored to detailed and complex scientific figure captions: a modified version of FActScore (Min et al., 2023), and G-Eval (Liu et al., 2023b). The modified FActScore breaks down the generated caption $y$ into a set of atomic units, denoted as $\mathcal{A}{y}$. Each atomic unit represents an independent description of either the overall figure or individual sub-figures. We then evaluate whether each atomic unit is supported by the ground-truth caption $\mathcal{C}$. For fine-grained evaluation, the LLM assigns a score $\phi(a,\mathcal{C})$ to each atomic unit $a\in\mathcal{A}{y}$ on a scale from 0 to 1, representing the degree of support from the ground-truth caption. A brevity penalty is incorporated to account for overly concise captions. The overall formulation is defined as follows:
$\begin{gathered}f(y)=\frac{1}{|\mathcal{A}y|}\sum{a\in\mathcal{A}_{y}}\phi(a,\mathcal{C})\cdot\text{exp}(\text{min}(1-\frac{\gamma}{\mathcal{A}_{y}},0)).\end{gathered}$
@@
We set $\gamma$ to 10 in our evaluation. This metric focuses on precision rather than recall. The second metric, G-Eval, compares the generated caption with the reference caption on a scale of 1 to 5, focusing on overall quality.
@@
MMSciQA: Figure Caption Matching.  We construct multiple-choice questions to evaluate models’ ability to match figures with their correct captions across three settings: (1) Figure-to-Caption (Fig2Cap): Models should select the correct main caption from four options, where distractors are drawn from other figures within the same article. This setting tests holistic figure comprehension. (2) Subfigure-to-Subcaption (SubFig2Cap): Given a random sub-figure, models must identify its corresponding sub-caption among four choices drawn from the same figure. This evaluates the ability to interpret specific components within a complex figure. (3) Subcaption-to-Subfigure (SubCap2Fig): In this reverse setting, given a sub-caption, models must select its matching sub-figure from all sub-figures within the same figure. This tests the model’s ability to associate textual descriptions with specific visual elements.
Data Split.  We allocated 1% of articles from each subject to both test and validation sets, yielding 1,418 test and 1,414 validation articles (5-50 articles per subject). Test samples were derived from unique articles to prevent content overlap. For caption tasks, we required a minimum length of 50 words. Each task setting comprised approximately 1,200 samples, balancing coverage and evaluation costs.